OpenStudy (unofficialllyy):

Can someone please help me with these 3 geometry questions? WILL FAN AND MEDAL!

OpenStudy (unofficialllyy):

The theme park company is building a scale model of the killer whale stadium main show tank for an investor's presentation. Each dimension will be made 6 times smaller to accommodate the mock-up in the presentation room. How many times smaller than the actual volume is the volume of the mock-up?

OpenStudy (unofficialllyy):

@phi

OpenStudy (unofficialllyy):

@Vocaloid

Vocaloid (vocaloid):

consider the relationship between dimensions and volume new length = (1/6)(old length) new width = (1/6)(old width) new depth = (1/6)(old depth) new volume = (1/6)*(1/6)*(1/6)*(old volume)

OpenStudy (unofficialllyy):

So it's basically \[\frac{ 1 }{ 6 }^3*(359,189)\] 359,189 was the volume for the main show tank in the first part of the problem that I did.

Vocaloid (vocaloid):

well, yeah, but it's not exactly what the question is asking it just wants you say compare the mockup volume and the actual volume "the mock-up volume is 6^3 times smaller than the actual volume" should suffice as an answer

OpenStudy (unofficialllyy):

Oh I see. Okay here's the next part Using the information from #4, answer the following question by filling in the blank: The volume of the actual tank is ____% of the mock-up of the tank

OpenStudy (unofficialllyy):

@sourwing

OpenStudy (unofficialllyy):

@nincompoop

OpenStudy (unofficialllyy):

@eyust707

OpenStudy (unofficialllyy):

@jim_thompson5910 when you are online, could you check this out? :)

Vocaloid (vocaloid):

"The volume of the actual tank is ____% of the mock-up of the tank" means 100*(actual volume)/(mock-up volume)

OpenStudy (unofficialllyy):

So I need to find out the mock up volume. How would i do that?

Vocaloid (vocaloid):

mock-up volume = (1/6)^3 * actual volume

OpenStudy (unofficialllyy):

So like i had before so that would be 1662.91203704 ?

Vocaloid (vocaloid):

yeah

OpenStudy (unofficialllyy):

Then I would do \[100*\frac{ 359189 }{1662.91203704 }?\]

OpenStudy (unofficialllyy):

Which is 21599.99999996

OpenStudy (unofficialllyy):

@Vocaloid

Vocaloid (vocaloid):

right, just make sure to round and add a percent sign at the end

OpenStudy (unofficialllyy):

So that rounded would be 216%? Im not very good with rounding

OpenStudy (unofficialllyy):

@Vocaloid

Vocaloid (vocaloid):

21600%

OpenStudy (unofficialllyy):

Okay this is the last part If you were to take a cross section parallel to the base of one of the holding tanks, how would you describe the shape?

Vocaloid (vocaloid):

that's kind of a vague question, but the cross-section would be a rectangle, assuming we'r e just looking at the holding tank part

OpenStudy (unofficialllyy):

What is a cross section?

Vocaloid (vocaloid):

uh, it's a bit tough to explain but I'll try